OpenStudy (anonymous):

what is the equation of a line that passes through the point (-2,1)?

OpenStudy (anonymous):

you cant get a line equation only with one point

myininaya (myininaya):

y=mx+b is a line we aren't given any info about the slope so i guess we can use any slope we want but it does the equation not a equation so m is the slope we know a point on the line so we can find b (the yintercept) 1=-2m+b 1+2m=b 1+2m is the y-intercept so the equation of the line is y=mx+(1+2m)

OpenStudy (anonymous):

you need another point or a slope

OpenStudy (anonymous):

the multiple choice answers for this problem are A) y=-x + 1 B) y=-x -2 C) y=1-x Dy=2-x so any ideas on how to get on of the answers? by the way myinaya you are good at explaining things.

myininaya (myininaya):

so i guess it is looking for a equation so we have y=mx+(1+2m) lets see which of the above lines fit this they all have slope -1 right? m=-1 then so y=-1x+(1+2(-1)) y=-x+(1-2) y=-x-1

myininaya (myininaya):

i don't see this choice

OpenStudy (anonymous):

oh yah that is a choice actually. sorry i wrote it wrong!

myininaya (myininaya):

good lol i was fixing to say none of these lol

myininaya (myininaya):

i go to bed gl with your math problems :)

OpenStudy (anonymous):

wait i have a question... how ddo you know that they have a -1 slope

myininaya (myininaya):

one more thing you could went through all of them in plug in x=-2 and y=1 and if you got the same thing on both sides then that is the line you are looking for like for instance A is y=-x+1 A is wrong since 1 does not equal 3 B is y=-x-2 B is wrong since 1 does not equal 0 C is y=1-x C is wrong since 1 does not equal 3 D is y=2-x D is wrong since 1 does not equal 4 also A and C are the same equation i bet it was one those you wrote wrong

myininaya (myininaya):

all them are pretty much in y=mx+b form the coeifficient of x is the slope they all have -1 in front of x we can only determine the slope if the line is written in y=mx+b form by the way not ax+by=c it has to be y=mx+b

myininaya (myininaya):

y=-x-1 is the line we found to be the right one (-2,1) x=-2 and y=1 1=-(-2)-1=1 since we have the same thing on both sides this is the line that (-2,1) is on

OpenStudy (anonymous):

one last question and then you can go to bed.... How do you solve ....what is the equation of the line that passes through points (7,-4) and (-3,5)?

myininaya (myininaya):

a line has the form y=mx+b so we need to know two things: slope=m and the y-intercept=b so we can find the slope because we are given two points m=slope=change of y/change of x=(5-(-4))/(-3-7)=9/(-10)=-9/10 so we know m so can fill one blank y=-9/10*x+b we know a point on the line (you can use either one-you will get the same thing) say we choose (-3,5) x=-3 y=5 so we only have one unknown in the equation y=-9/10*x+b since we know a point so we have 5=-9/10*(-3)+b 5=27/10+b 5-27/10=b 50/10-27/10=b (found common denominator) 23/10=b so we know both m and b now so we can now write our equation y=-9/10*x+23/10

myininaya (myininaya):

you can also write the line in general form ax+by=c multiply 10 on bot sides 10y=-9x+23 add 9x on both sides 10y+9x=23 9x+10y=23

myininaya (myininaya):

good?

myininaya (myininaya):

goodnight?

OpenStudy (anonymous):

sory my computer froze... the answer choices are y+4=-9/10(x-7) or y-4 = -9/10(x + 7) which one?

myininaya (myininaya):

ok so we see the form it is written in y-y0=m(x-x0) (x0,y0)=(7,-4) m=-9/10 so y-(-4)=-9/10(x-7) y+4=-9/10(x-7)

myininaya (myininaya):

the first one

OpenStudy (anonymous):

thx.goodnight

myininaya (myininaya):

(x0,y0) is a point on the line and (7,-4) was the point they used so thats why i chose that point instead of the other but it doesn't matter what point you do use because you can manipulate both to show that they are the same

myininaya (myininaya):

goodnight :)